OpenStudy (anonymous):

A skyscraper casts a shadow 200 feet long. If the angle of elevation of the sun is 49 , then the height of the skyscraper is _____.

OpenStudy (mathstudent55):



OpenStudy (anonymous):

tan(49)=x/200

OpenStudy (anonymous):

or in the dude above me's drawing, tan(49)=y/200

OpenStudy (anonymous):

Im not really good at tan can you elaborate

OpenStudy (mathstudent55):

In the above picture, 200 ft is the adjacent leg to the 49-deg angle. y is the opposite leg to teh 49-deg angle. Which trig function (sine, cosine, or tangent) relates the adjacent to the opposite legs?

OpenStudy (anonymous):

just use a calculator online. tan(49) is some random number BLAH. so BLAH = y/200 -> 200(BLAH) = y where y is your answer

OpenStudy (anonymous):

and yea tan stands for tangent, cos for cosine, and sin for sine

OpenStudy (anonymous):

Tan

OpenStudy (mathstudent55):

Great. \(\tan 49^o = \dfrac{\text{opp}}{\text{adj}} \) \( \tan 49^o = \dfrac{y}{200 ~ft} \) Is this ok so far?

OpenStudy (anonymous):

Yes that good so far....when i found tan 49 does it equal to -3.1729... Or something else cause mi phone calculator isnt advance

OpenStudy (mathstudent55):

First, your calculator is in radians mode. Set it to degree mode.

OpenStudy (mathstudent55):

tan 49 rad = -3.1729 tan 49 deg = 1.1504

OpenStudy (mathstudent55):

\( 1.1504 = \dfrac{y}{200 ~ft} \) Multiply both sides by 200 ft. \(y = 1.1504 \times 200 ~ft \)

OpenStudy (anonymous):

Oh okay thank you so much

OpenStudy (mathstudent55):

wlcm